OpenStudy (anonymous):

Uranium-232 has a half-life of 72 years. How many years will it take for a sample of uranium-232 to decay to 80% of its initial quantity?

OpenStudy (anonymous):

\[A_0e^{rt}=A\] \[e^{r(72)}=1/2\] \[ln(e^{72r})=ln(1/2)\] \[72rln(e)=ln(1/2)\] \[72r=ln(1/2)\]

OpenStudy (anonymous):

Thanks :/